OpenStudy (anonymous):

If f(x)=F^-1(x) then f(x) may equal:

OpenStudy (anonymous):

is there another function for which \( f(x)=f^{-1} (x) \) ?

OpenStudy (anonymous):

@mukushla I think this isn't the full question ! There is a lot of function that satify f^-1(x)=f(x)

OpenStudy (anonymous):

i dont know we must prove it...

OpenStudy (anonymous):

for example 1/x !

OpenStudy (anonymous):

yes u r right @Neemo

OpenStudy (anonymous):

f(x)=f −1 (x) the f(x) may equal: A) 2^x B) 1/2x c) 1 D) 1/x E 1/SQ rt of X are the options

OpenStudy (anonymous):

It must be strictly increasing...then we can say that is f(x)=x

OpenStudy (anonymous):

the answer is 1/ (1/x)=x just don't know how to get to that point